# The neutrino luminosity and energy spectrum of nova outburst  [^1] Released on March, 1st, 2021

### v1

@@ Abstract
The nova outburst can produce a large number of neutrinos, whether it is the nuclear reaction process during the explosion or the shock wave acceleration proton process. We study the low-energy nuclear and thermal neutrino luminosity of novae with CO white dwarf (WD) mass ranging from 0.6 to 1.1 $\rm M_{\odot}$ with different accretion rates $\dot{M}$, core temperatures $(T_{\mathrm{C}})$, and mixing degrees. We find that during the accretion phase, low-energy neutrinos are mainly produced by pp chains and plasma decay, and photon luminosity is greater than low-energy nuclear and thermal neutrino luminosity. During the thermonuclear runaway (TNR) phase, low-energy neutrinos are mainly produced by the CNO cycle and photon-neutrino, and the low-energy nuclear and thermal neutrino luminosity far exceeds the photon luminosity. We find that the more massive the WD, the shorter the cycle time and the higher the low-energy nuclear neutrino luminosity. The higher the accretion rate, the lower the low-energy nuclear neutrino luminosity. If the accretion mixing effect is not taken into account, the outburst interval becomes longer, the low-energy nuclear neutrino luminosity will be increased. And for the cooler nova model $(T_{\mathrm{C}}=1\times 10^{7}\rm K)$, the low-energy nuclear neutrino luminosity will be lower during the accretion phase and higher at the TNR. We also predict the neutrino luminosity and energy spectrum of the upcoming recurrent nova T Coronae Borealis (T CrB). We estimate that the next T CrB outburst has a low-energy nuclear neutrino peak luminosity of $2.7\times 10^{8}\ \rm L_{\nu,\odot}$ and a low-energy nuclear neutrino outburst duration of 88 days. In addition, we predict that the high-energy hadronic neutrino flux produced by T CrB nova can not be observed by the current-generation IceCube.
@@ 1 Introduction
Neutrinos are considered the most abundant particles in the universe and play a crucial role in the evolution of stars(Pandey, 2024). For example, helium flashes (Catelan et al., 1996), thermal pulses (Shi et al., 2020), and cooling of WD (Fontaine et al., 2001) in low-mass stars, late stage evolution of high-mass stars (Woosley et al., 2002), supernova explosion (Janka, 2017), the cooling of neutron stars (Nomoto and Tsuruta, 1981). In addition, the study of neutrinos has significantly contributed to our comprehension of the fundamental properties and of the Universe. They can probe neutrino interactions at high energies, and explore deviations from the Standard Model of particle physics. For example, measuring the cross section of neutrino and nucleon interaction for energies larger than $\rm 1\ TeV$ (Reno, 2023) and prove that Leptoquarks can change the neutrino cross-section (Klein, 2019). High-energy neutrinos are only produced at the same time as high-energy gamma rays (gamma-ray) through the hadron process, so detecting high-energy neutrinos can also determine the origin of high-energy gamma-ray (Song, 2023). Meanwhile, the flavour composition of high-energy neutrinos can be a probe of neutrino magnetic moments (Popov and Studenikin, 2024). Neutrinos carry invaluable information about the existence (or absence) of energetic protons and shed light on the location of the $\gamma$-ray production region (Sahakyan et al., 2014).
@@
Neutrino detection is rapidly developing as a new astronomical messenger. Super Kamiokande (SK) is a detector used to study neutrino Cherenkov radiation from different sources(Fukuda et al., 2003). It primarily detects and analyzes particle interactions within the energy range of several MeV to a few tens of TeV, such as solar neutrino(Abe et al., 2016) and supernova neutrino(Mori et al., 2022). The IceCube detector, the first of its kind with a gigaton-scale capacity, was primarily designed to observe neutrinos originating from the most violent astrophysical phenomena in our universe (IceCube Collaboration et al., 2006). Such as blazar (Collaboration et al., 2018; Tavecchio, 2018; Oikonomou et al., 2019; Cao et al., 2024), gamma-ray bursts (Biehl et al., 2018), black holes (Ma and Wang, 2024; van Velzen et al., 2024), active galactic (Kurahashi et al., 2022) and tidal disruption events (Winter and Lunardini, 2021). These neutrinos usually have extremely high energies, ranging from TeV to PeV and beyond.
@@
Novae are significant sites for the production of neutrinos (Razzaque et al., 2010). The novae are usually binary systems (Warner, 1995), and the primary star is a white dwarf (WD). According to the type of companion star, it can be divided into two categories: one is the classical nova, the companion star is the main sequence star, and the other is the symbiotic nova, the companion star is the red giant (RG). When a sufficient number of layers of material have accumulated, it eventually triggers a thermonuclear runaway (TNR) explosion on the surface of the WD (Hernanz, 2005), and significantly enhances the luminosity of the WD to approximately $\rm 10^{4-5}\ L_{\odot}$ (Gomez-Gomar et al., 1998; Hellier, 2001; Knigge et al., 2011), such a violent nuclear reaction would produce a large number of electron neutrinos. Meanwhile, the shock wave from the nova outburst accelerates the surrounding particles to produce neutrinos. The shock wave can be categorized into two distinct types: internal shock waves and external shock waves. Internal shock wave can be attributed to the collision between slower-moving ejecta and faster-moving ejecta (Metzger et al., 2015). The fast-moving outflow has the potential to collide with a pre-existing dense wind emanating from the RG star, resulting in the generation of an external shock (Lü et al., 2011; Martin and Dubus, 2013). Symbiotics are therefore more likely to produce meaningful hadronic emission at their forward shocks. However, classical novae can and do produce hadronic emission at internal shocks (Metzger et al., 2015; Ackermann et al., 2014; Cheung et al., 2016; Franckowiak et al., 2018). In the “hadronic” scenario, the particles accelerated by the shock wave undergo interactions with the surrounding matter (proton-proton (pp) or proton-gamma ($\rm p\gamma$) collisions), resulting in the generation of neutral and charged muons. These muons subsequently decay into high-energy $(\rm>1\ GeV)$ neutrinos and gamma rays (Abbasi et al., 2022). The maximum energies of high-energy particles will be contingent upon the duration of the nova, the efficiency of the acceleration mechanism, and the energy losses incurred during cooling (Acciari et al., 2022). The particles are accelerated by external shock waves to produce high-energy neutrinos, which typically occur in symbiotic novae (De Sarkar et al., 2023).
@@
T Coronae Borealis (T CrB) is a classical symbiotic recurrent nova (Iłkiewicz et al., 2023). The primary star of this type of nova is in close proximity to the Chandrasekhar limit, and it must undergo accretion at exceedingly high rates, approximately $1.7\times 10\ \rm M_{\odot}\ yr^{-1}$, in order to rapidly accumulate critical-mass envelopes and consequently outburst with such frequency (Kato, 1991). In addition to T CrB, there are three with red giant donors in the 10 known recurrent novae of the Galactic: V745 Sco, RS Ophiuchus, and V3890 Sgr. The WD mass in T CrB is $1.2\pm 0.2\ \rm M_{\odot}$, the companion red giant has a mass of the red giant of $1.12\pm 0.23\ \rm M_{\odot}$. The average accretion rate $(\dot{M}_{\mathrm{a}})$ during the superactive state is $2\times 10^{-8}\ \rm M_{\odot}\ yr^{-1}$ and a maximum of $4\times 10^{-8}\ \rm M_{\odot}\ yr^{-1}$ (Zamanov et al., 2023). The distance from T CrB to Earth is approximately 914 pc. And many groups predict it will outburst in $2024$ (Schaefer, 2023; Maslennikova et al., 2023). Prior to this, the T CrB is recorded eruptions in 1217.8, 1789.9, 1866.4 and 1946.1, the time span between bursts is approximately 80 years (Schaefer, 2023). Because T CrB is very close to us and is about to outburst, it is necessary to study its neutrino luminosity and energy spectrum.
@@
In this paper, we plan to investigate neutrino luminosity and energy spectrum during nova outburst. Section 2 gives the details of software instruments, input physics, reaction networks and model selection. In section 3 we discuss the low-energy nuclear and thermal neutrino luminosity during nova outburst and the influence of different $M_{\rm{WD}}$, $\dot{M}$ and $T_{\mathrm{C}}$ on low-energy nuclear neutrino luminosity. In section 4 we predict the neutrino energy spectrum when the T CrB next outbursts. The conclusions are given in section 5.
2 Nova MODEL
@@
Classical novae eruptions are the result of unstable hydrogen burning on the surface of either CO WD or ONeMg WD cores, which accrete hydrogen-rich materials from their main sequence (MS) or red giant (RG) companions in low-mass, close binary systems. We employ the open-source stellar evolution code (MESA, version 23.05, Paxton et al. 2011, 2013, 2015, 2018; Paxton 2019; Jermyn et al. 2023) to construct nova models. The MESA is capable of generating state-of-the-art simulations of novae (Denissenkov et al., 2013; Zhu et al., 2021).
@@
The neutrino normalization $L_{\nu,\odot}=0.02398\cdot L_{\gamma,\odot}=9.1795\times 10^{31}\mathrm{erg\ s}^{-1}$ is adopted in Farag et al. (2020).
2.1 Neutrinos produced by nuclear reaction and thermal processes
The generation of neutrinos inside stars can be broadly categorized into two distinct processes. One is weak nuclear processes, including $\beta^{\pm}$ decay and $e^{\pm}$ capture (Kato et al., 2020):
@@ 2.1 Neutrinos produced by nuclear reaction and thermal processes
Such as $\rm p(p,e^{+}\nu)d$, $\rm p(e^{-}p,\nu_{e})d$, $\rm{}^{3}He(p,e^{+}\nu_{e})\alpha$ in pp chains, $\rm{}^{13}N(,e^{+}\nu_{e})^{13}C$, $\rm{}^{15}O(,e^{+}\nu_{e})^{15}N$, $\rm{}^{17}F(,e^{+}\nu_{e})^{17}O$ in CNO cycle. For the nova model, we choose pp-and-cno-extras.net as a network of nuclear reactions during nova outburst. The occurrence of classical novae eruptions can be attributed to the instability in hydrogen burning on the surface of WD cores. Therefore, the most important nuclear reactions during nova eruption are the pp chain and the CNO cycle (Denissenkov et al., 2013).
The other is thermal processes, which are sensitive to temperature and density. Thermal neutrino energy losses are from Itoh et al. (1996), including:
@@
photon-neutrino$\rm\left(T<4\times 10^{8}\ K,\rho<10^{5}\ g\ cm^{-3}\right)$
@@
bremsstrahlung$\rm\left(10^{8}<\rho<10^{10}\ g\ cm^{-3}\right)$
@@
The MESA nuclear reaction rates are a combination of rates from NACRE (Angulo et al., 1999) and JINA REACLIB Cyburt et al. (2010). The treatment of screening corrections, as proposed by (Chugunov et al., 2007), serves to enhance nuclear reaction rates in dense plasmas. The weak reaction rates utilized in this study are derived from previous works (Langanke and Martínez-Pinedo, 2000; Oda et al., 1994; Fuller et al., 1985).
2.2 Neutrinos produced by $pp$ collisions
T CrB is a recurrent nova, the ambient environments and explosion mechanisms are similar to RS Oph (Evans et al., 2019), when the ejection material encounters the wind material around the red giant, it will cause a shock system (Zamanov et al., 2023).
The interaction between RG wind and the ejecta will produce two shocks, a forward shock propagating into the RG wind, and a reverse shock propagating back into the ejecta (Zheng et al., 2022). According to Aharonian et al. (2022), we also only consider the forward shock for explaining the neutrino emission.
In the early stage, the ejecta expands freely and is almost unaffected by the interstellar medium surrounding the binary. We assume when the distance between the wind and the nova r is much larger than the semi-major axis of the nova $a$, the structure of the wind can be approximately spherical, and the ejecta is assumed to be concentrated at the front of the shock wave with a thickness of $hr_{\mathrm{sh}}$, h is $1/10$ of the shock radius. Meanwhile, we are only interested in the initial day of the T CrB outburst, thus we do not consider the radiative cooling phase after the shock wave deceleration. We assume that the shock radius $(r_{\mathrm{sh}})$ and velocity $(v_{\mathrm{sh}})$ of the T CrB are $\rm 4\times 10^{13}\ cm$ and $\rm 4500\ km/s$ (Zheng et al., 2024), respectively. The T CrB companion has a mass of $1\ \rm M_{\odot}$ (Pavlenko et al., 2020) and a typical mass loss rates $(\dot{M}_{\mathrm{RG}})$ of $5\times 10^{-7}\ \rm M_{\odot}\ \mathrm{yr}^{-1}$ (Ferrarotti and Gail, 2006). The velocity of RG wind $(v_{\mathrm{RG}})$ is $\rm 10\ km/s$.
The accelerated protons in the forward shock will interact with the matter in the RG wind, mainly the collision of protons and protons, and the resulting charged pions and kaons decay produces observable gamma-ray and high-energy neutrinos. The density of the matter in the RG wind can be estimated as:
@@ 2.2 Neutrinos produced by $pp$ collisions
Neutrino energy spectrum is also sensitive to the maximum accelerated proton energy, and it is to relate the acceleration timescale with the age of the system. If particles are accelerated via diffusive shock acceleration, then the acceleration time should be of order the diffusion time ($D/v_{\mathrm{sh}}^{2}$). For Bohm diffusion, the characteristic time scale of particle acceleration can be estimated as:
@@
where $\xi\leq(v_{sh}/c)^{2}\approx 2\times 10^{-4}$ is acceleration parameter, $R_{\mathrm{L}}=E/eB$ is the Larmor radius of particles in the magnetic field B (in Gauss). According to Bednarek and Pabich (2011) and Marcowith et al. (2018), the proton Larmor radius for a particle energy of 1 GeV can be written as :
@@
As discussed in Chomiuk et al. (2012), the magnetic field B is
@@
where $T_{\mathrm{RG}}\rm$ is the temperature of the RG wind, $k_{B}=1.380649\times 10^{-23}\ \rm J/K$. Based on Stanishev et al. (2004), the RG companion star of T CrB has an effective surface temperature of 3500 K. However, the $T_{\mathrm{RG}}\rm$ produced from the companion star should be less than the surface temperature (Chomiuk et al., 2012). Following De Sarkar et al. (2023) for nova RS Oph, we set $T_{\mathrm{RG}}=1000\rm\ K$. The primary mechanism for hadron energy loss occurs through collisions with the matter present in RG winds. An estimation of the timescale for energy losses due to pion production in $pp$ collisions can be derived from (Diesing et al., 2023):
@@
where $\sigma_{pp}\approx 3\times 10^{-26}\rm\ cm^{2}$ is the total cross section of $pp$ interactions, $k=0.5$ is the inelasticity coefficient in this collision. By comparing the time scale for the energy losses in $pp$ collisions and the characteristic acceleration time scale ($\tau_{\mathrm{pp}}=\tau_{\mathrm{acc}}$), the maximum proton energy can be estimated as :
@@
According to Caprioli (2023), the diffusion length ( $D/v_{\mathrm{sh}}$) of the particles with energy $E_{max}$ can not to exceed the source size $r_{\mathrm{sh}}$, which is similar with Hillas criterion. For our hadronic model, we find that the accelerated proton population can reach a maximum energy of approximately 330 GeV. As described by (Caprioli et al., 2010), the maximum energy is usually the location of the cut-off energies, so we set $E_{0}=330\rm\ GeV$. The results obtained from the nova model discussed in this section will be presented in the subsequent section.
3 Result

### v2

# Neutrino luminosity and the energy spectrum of a nova outburst  [^1] Released on March, 1st, 2021

@@ Abstract
The nova outburst can produce a large number of neutrinos, whether it is the nuclear reaction process during the explosion or the shock wave acceleration proton process. We study the low-energy nuclear and thermal neutrino luminosity of novae with CO white dwarf (WD) mass ranging from 0.6 to 1.1 $\rm M_{\odot}$ with different accretion rates $\dot{M}$, core temperatures $(T_{\mathrm{C}})$, and mixing degrees. We find that during the accretion phase, low-energy neutrinos are mainly produced by pp chains and plasma decay, and photon luminosity is greater than low-energy nuclear and thermal neutrino luminosity. During the thermonuclear runaway (TNR) phase, low-energy neutrinos are mainly produced by the CNO cycle and photon-neutrino interaction, and the low-energy nuclear and thermal neutrino luminosity far exceeds the photon luminosity. We find that the more massive the WD, the shorter the cycle time and the higher the low-energy nuclear neutrino luminosity. The higher the accretion rate, the lower the low-energy nuclear neutrino luminosity. If the accretion mixing effect is not taken into account, the outburst interval becomes longer, and the low-energy nuclear neutrino luminosity will be increased. For the cooler nova model $(T_{\mathrm{C}}=1\times 10^{7}\rm K)$, the low-energy nuclear neutrino luminosity will be lower during the accretion phase and higher at the TNR. We also predict the neutrino luminosity and energy spectrum of the upcoming recurrent nova T Coronae Borealis (T CrB). We estimate that the next T CrB outburst has a low-energy nuclear neutrino peak luminosity of $2.7\times 10^{8}\ \rm L_{\nu,\odot}$ and a low-energy nuclear neutrino outburst duration of 88 days. In addition, we predict that the high-energy hadronic neutrino flux produced by T CrB nova cannot be observed by the current-generation IceCube.
@@ I Introduction
Neutrinos are considered the most abundant particles in the Universe and play a crucial role in the evolution of stars[1], for example, helium flashes [2], thermal pulses [3], and cooling of white dwarfs (WDs) [4] in low-mass stars, late stage evolution of high-mass stars [5], supernova explosion [6], and the cooling of neutron stars [7]. In addition, the study of neutrinos has significantly contributed to our comprehension of the fundamental properties of the Universe. It can probe neutrino interactions at high energies and explore deviations from the Standard Model of particle physics, for example, measuring the cross section of neutrino and nucleon interaction for energies larger than $\rm 1\ TeV$ [8] and proving that leptoquarks can change the neutrino cross section [9]. High-energy neutrinos are only produced at the same time as high-energy $\gamma$ rays through the hadron process, so detecting high-energy neutrinos can also determine the origin of high-energy $\gamma$ rays [10]. Meanwhile, the flavor composition of high-energy neutrinos can be a probe of neutrino magnetic moments [11]. Neutrinos carry invaluable information about the existence (or absence) of energetic protons and shed light on the location of the $\gamma$-ray production region [12].
@@
Neutrino detection is rapidly developing as a new astronomical messenger. Super-Kamiokande (SK) is a detector used to study neutrino Cherenkov radiation from different sources[13]. It primarily detects and analyzes particle interactions within the energy range of several MeV to a few tens of TeV, such as solar neutrinos[14] and supernova neutrinos[15]. The IceCube detector, the first of its kind with a gigaton-scale capacity, was primarily designed to observe neutrinos originating from the most violent astrophysical phenomena in our Universe [16], such as blazars [17, 18, 19, 20], $\gamma$-ray bursts [21], black holes [22, 23], active galactic nuclei, [24] and tidal disruption events [25]. These neutrinos usually have extremely high energies, ranging from TeV to PeV and beyond.
@@
Novae are significant sites for the production of neutrinos [26]. The novae are usually binary systems [27], and the primary star is a white dwarf. According to the type of companion star, it can be divided into two categories: one is the classical nova, where the companion star is the main sequence star, and the other is the symbiotic nova, where the companion star is the red giant (RG). When a sufficient number of layers of material have accumulated, it eventually triggers a thermonuclear runaway (TNR) explosion on the surface of the WD [28] and significantly enhances the luminosity of the WD to approximately $\rm 10^{4-5}\ L_{\odot}$ [29, 30, 31]; such a violent nuclear reaction would produce a large number of electron neutrinos. Meanwhile, the shock wave from the nova outburst accelerates the surrounding particles to produce neutrinos. The shock wave can be categorized into two distinct types: internal and external shock waves. Internal shock waves can be attributed to the collision between slower-moving ejecta and faster-moving ejecta [32]. The fast-moving outflow has the potential to collide with a preexisting dense wind emanating from the RG star, resulting in the generation of an external shock [33, 34]. Symbiotics are therefore more likely to produce meaningful hadronic emission at their forward shocks. However, classical novae can and do produce hadronic emission at internal shocks [32, 35, 36, 37]. In the “hadronic” scenario, the particles accelerated by the shock wave undergo interactions with the surrounding matter [proton-proton (pp) or proton-$\gamma$ ($\rm p\gamma$) collisions], resulting in the generation of neutral and charged muons. These muons subsequently decay into high-energy $(\rm>1\ GeV)$ neutrinos and $\gamma$ rays [38]. The maximum energies of high-energy particles will be contingent upon the duration of the nova, the efficiency of the acceleration mechanism, and the energy losses incurred during cooling [39]. The particles are accelerated by external shock waves to produce high-energy neutrinos, which typically occur in symbiotic novae [40].
@@
T Coronae Borealis (T CrB) is a classical symbiotic recurrent nova [41]. The primary star of this type of nova is in close proximity to the Chandrasekhar limit, and it must undergo accretion at exceedingly high rates, approximately $1.7\times 10^{-8}\ \rm M_{\odot}\ yr^{-1}$, in order to rapidly accumulate critical-mass envelopes and consequently outburst with such frequency [42]. In addition to T CrB, there are three with red giant donors in the ten known recurrent novae of the Galaxy: V745 Sco, RS Ophiuchus, and V3890 Sgr. The WD mass in T CrB is $1.2\pm 0.2\ \rm M_{\odot}$, and the companion red giant has a mass of $1.12\pm 0.23\ \rm M_{\odot}$. The average accretion rate $(\dot{M}_{\mathrm{a}})$ during the superactive state is $2\times 10^{-8}\ \rm M_{\odot}\ yr^{-1}$ with a maximum of $4\times 10^{-8}\ \rm M_{\odot}\ yr^{-1}$ [43]. The distance from T CrB to Earth is approximately 914 pc. Many groups predicted it will outburst in 2024 [44, 45]. Prior to this, the T CrB is recorded eruptions in 1217.8, 1789.9, 1866.4, and 1946.1; the time span between bursts is approximately 80 yr [44]. Because T CrB is very close to us and is about to outburst, it is necessary to study its neutrino luminosity and energy spectrum.
@@
In this paper, we plan to investigate neutrino luminosity and energy spectrum during nova outburst. Sec. II gives the details of software instruments, input physics, reaction networks, and model selection. In Sec. III we discuss the low-energy nuclear and thermal neutrino luminosity during nova outburst and the influence of different $M_{\rm{WD}}$, $\dot{M}$, and $T_{\mathrm{C}}$ on low-energy nuclear neutrino luminosity, we also predict the neutrino energy spectrum when the T CrB next outbursts. The conclusions are given in Sec. IV.
II Nova MODEL
@@
Classical novae eruptions are the result of unstable hydrogen burning on the surface of either CO WD or ONeMg WD cores, which accrete hydrogen-rich materials from their main sequence or RG companions in low-mass, close binary systems. We employ the open-source stellar evolution code mesa, version 23.05 (Paxton $et\ al$. [46, 47, 48, 49, 50] and Jermyn et al. [51]) to construct nova models. mesa is capable of generating state-of-the-art simulations of novae [52, 53].
@@
The neutrino normalization $L_{\nu,\odot}=0.02398\cdot L_{\gamma,\odot}=9.1795\times 10^{31}\ \mathrm{erg\ s}^{-1}$ is adopted in Farag et al. [67].
II.1 Neutrinos produced by nuclear reaction and thermal processes
The generation of neutrinos inside stars can be broadly categorized into two distinct processes. The first includes weak nuclear processes, including $\beta^{\pm}$ decay, electron capture (EC), and positron capture (PC) [68],
@@ II.1 Neutrinos produced by nuclear reaction and thermal processes
such as $\rm p(p,e^{+}\nu)d$, $\rm p(e^{-}p,\nu_{e})d$, and $\rm{}^{3}He(p,e^{+}\nu_{e})\alpha$ in pp chains and $\rm{}^{13}N(,e^{+}\nu_{e})^{13}C$, $\rm{}^{15}O(,e^{+}\nu_{e})^{15}N$, and $\rm{}^{17}F(,e^{+}\nu_{e})^{17}O$ in the CNO cycle. For the nova model, we chose pp-and-cno-extras.net as a network of nuclear reactions during nova outburst. The occurrence of classical novae eruptions can be attributed to the instability in hydrogen burning on the surface of WD cores. Therefore, the most important nuclear reactions during nova eruption are the pp chain and the CNO cycle [52].
The other includes thermal processes, which is sensitive to temperature and density. Thermal neutrino energy losses are from Itoh et al. [69], including the following:
@@
photon-neutrino interaction$\rm\left(T<4\times 10^{8}\ K,\rho<10^{5}\ g\ cm^{-3}\right)$
@@
and bremsstrahlung$\rm\left(10^{8}<\rho<10^{10}\ g\ cm^{-3}\right)$
@@
The mesa nuclear reaction rates are a combination of rates from NACRE [70] and JINA REACLIB [71] databases. The treatment of screening corrections, as proposed by Chugunov et al. [72], serves to enhance nuclear reaction rates in dense plasmas. The weak reaction rates utilized in this study are derived from previous works [73, 74, 75].
II.2 Neutrinos produced by $pp$ collisions
T CrB is a recurrent nova; the ambient environments and explosion mechanisms are similar to RS Ophiuchus (RS Oph) [76]. When the ejection material encounters the wind material around the red giant, it will cause a shock system [43].
The interaction between RG wind and the ejecta will produce two shocks, a forward shock propagating into the RG wind and a reverse shock propagating back into the ejecta [77]. According to Aharonian et al. [78], we also only consider the forward shock for explaining the neutrino emission.
In the early stage, the ejecta expands freely and is almost unaffected by the interstellar medium surrounding the binary. We assume when the distance between the wind and the nova $r$ is much larger than the semimajor axis of the nova $a$, the structure of the wind can be approximately spherical, and the ejecta is assumed to be concentrated at the front of the shock wave with a thickness of $hr_{\mathrm{sh}}$, where $h$ is $1/10$ of the shock radius. Meanwhile, we are only interested in the initial day of the T CrB outburst, thus we do not consider the radiative cooling phase after the shock wave deceleration. We assume that the shock radius $(r_{\mathrm{sh}})$ and velocity $(v_{\mathrm{sh}})$ of the T CrB are $\rm 4\times 10^{13}\ cm$ and $\rm 4500\ km/s$ [79], respectively. The T CrB companion has a mass of $1\ \rm M_{\odot}$ [80] and a typical mass loss rate $(\dot{M}_{\mathrm{RG}})$ of $5\times 10^{-7}\ \rm M_{\odot}\ \mathrm{yr}^{-1}$ [81]. The velocity of RG wind $(v_{\mathrm{RG}})$ is $\rm 10\ km/s$.
The accelerated protons in the forward shock will interact with the matter in the RG wind, mainly the collision of protons and protons, and the resulting charged pion and kaon decay produces observable $\gamma$-ray and high-energy neutrinos. The density of the matter in the RG wind can be estimated as
@@ II.2 Neutrinos produced by $pp$ collisions
The neutrino energy spectrum is also sensitive to the maximum accelerated proton energy, and it is to relate the acceleration timescale with the age of the system. If particles are accelerated via diffusive shock acceleration, then the acceleration time should be of order of the diffusion time ($D/v_{\mathrm{sh}}^{2}$). For Bohm diffusion, the characteristic timescale of particle acceleration can be estimated as
@@
where $\xi\leq(v_{sh}/c)^{2}\approx 2\times 10^{-4}$ is acceleration parameter, $R_{\mathrm{L}}=E/eB$ is the Larmor radius of particles in the magnetic field B (in Gauss). According to Bednarek and Pabich [82] and Marcowith et al. [83], the proton Larmor radius for a particle energy of 1 GeV can be written as
@@
As discussed in Chomiuk et al. [84], the magnetic field B is
@@
where $T_{\mathrm{RG}}\rm$ is the temperature of the RG wind, $k_{B}=1.380649\times 10^{-23}\ \rm J/K$. Based on Stanishev et al. [85], the RG companion star of T CrB has an effective surface temperature of 3500 K. However, the $T_{\mathrm{RG}}\rm$ produced from the companion star should be less than the surface temperature [84]. Following De Sarkar et al. [40] for nova RS Oph, we set $T_{\mathrm{RG}}=1000\rm\ K$. The primary mechanism for hadron energy loss occurs through collisions with the matter present in RG winds. An estimation of the timescale for energy losses due to pion production in $pp$ collisions can be derived from [86],
@@
where $\sigma_{pp}\approx 3\times 10^{-26}\rm\ cm^{2}$ is the total cross section of $pp$ interactions, and $k=0.5$ is the inelasticity coefficient in this collision. By comparing the timescale for the energy losses in $pp$ collisions and the characteristic acceleration timescale ($\tau_{\mathrm{pp}}=\tau_{\mathrm{acc}}$), the maximum proton energy can be estimated as
@@
According to Caprioli [87], the diffusion length ($D/v_{\mathrm{sh}}$) of the particles with energy $E_{max}$ cannot exceed the source size $r_{\mathrm{sh}}$, which is similar to the Hillas criterion. For our hadronic model, we find that the accelerated proton population can reach a maximum energy of approximately 330 GeV. As described by Caprioli et al. [88], the maximum energy is usually the location of the cutoff energies, so we set $E_{0}=330\rm\ GeV$. The results obtained from the nova model discussed in this section will be presented in the subsequent section.
III Result